OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)?

OpenStudy (anonymous):

D = sqrt( (-2-3)^2 + (1-7)^2 ) = sqrt(61)

OpenStudy (anonymous):

distance formula is: \[d=\sqrt{\Delta x^2+\Delta y^2}\]So,\[d=\sqrt{[3-(-2)]^2+(7-1)^2}\]This is,\[d=\sqrt{25+36}\]Or,\[\sqrt{61}\]

OpenStudy (anonymous):

@eseidl: On a side note, how do you get the root symbol? It's much more convenient to understand!

OpenStudy (anonymous):

type this \sqrt{61}

OpenStudy (anonymous):

but you have to enclose it in the display math tags which i cannot write here because it will just display the math

OpenStudy (anonymous):

if you want i can show you in chat but i cannot show you here

OpenStudy (anonymous):

I push the "equation" button (beside "attach file") when typing a reply...the root symbol is within.